# Faster and Memory-Efficient Training of Sequential Recommendation Models for Large Catalogs

### v4

@@ 4.3. Training Details
For all datasets, we trained models on a single Nvidia A100 GPU. Prior to our experiments, we tuned the number of self-attention blocks, attention heads, and hidden size. As a result, the model configuration was set to two self-attention blocks with two attention heads and a hidden size $D$ of $256$. The models were trained using the Adam optimizer with a learning rate of 1e31e-3. The training was performed with mixed-precision (float16) using the PyTorch Lightning trainer (micikevicius2017mixed).
@@ 5.1. Optimal Accuracy-Memory Trade-off
The first research question is how to determine the optimal configuration of training hyperparameters given the memory constraints and how specific memory-related hyperparameters affect performance. We investigate how batch size (bs), the number of negatives (ns), and sequence length (sl) affect the NDCG@10 metric of the SASRec model after training. We used a fixed grid of configurations, evaluating approximately $170$ points per dataset, resulting in a total of $997$ runs across all datasets. While the grid does not provide a fully uniform coverage of the hyperparameter space, the analysis can still be interpreted under the assumption of reasonably even exploration (scarlett2017lower; srinivas2010gaussian).
@@ 5.4.1. Comparison Sampling Strategies
Tables 6, 7 present data to compare metrics and computational resource usage for training with $\mathrm{CCE}^{-}$ using global uniform and popularity-based sampling strategies. In terms of $\mathrm{NDSG@10}$ metric, global uniform sampling consistently outperforms the popularity-based approach across all considered datasets. However, popularity-based sampling shows significant gains in diversity-related metrics for the Gowalla and Beauty datasets. Compared to popularity-based sampling, using global uniform sampling leads to a reduction in $\mathrm{Coverage@10}$ by 185.8% and 63.9 %, and a reduction in $\mathrm{Surprisal@10}$ by 40% and 19.8%. These results align with observations that popularity-weighted sampling enhances exposure to long-tail items, improving recommendation diversity and novelty (wang2024popularitylearning).
@@
Another approach to negative sampling is the Scalable Cross-Entropy (SCE) method  (mezentsev2024scalable). This method approximates the CE loss by employing a sampling procedure that computes logits only for elements with the largest absolute values, rather than for all items in the catalog. As a result, memory usage for the logit tensor is significantly reduced, and the loss computation is accelerated.
@@ 5.5. Comparison of CCE and $\mathrm{CCE}^{-}$ with BCE
As outlined in the original paper (kang2018self), the SASRec model can be trained using the Binary Cross-Entropy (BCE) loss. To evaluate the effectiveness of our proposed methods, we conducted training SASRec with the BCE loss.
@@ 5.6. Analysis of efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec
In addition to the SASRec architecture, we study the efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec  (sun2019bert4rec).

### v5

@@ 4.3. Training Details
For all datasets, we trained models on a single Nvidia A100 GPU. Prior to our experiments, we tuned the number of self-attention blocks, attention heads, and hidden size. As a result, the model configuration was set to two self-attention blocks with two attention heads and a hidden size $D$ of $256$. The models were trained using the Adam optimizer with a learning rate of 1e31e-3. The training was performed with mixed-precision (float16) using the PyTorch Lightning trainer (Micikevicius et al., 2017).
@@ 5.1. Optimal Accuracy-Memory Trade-off
The first research question is how to determine the optimal configuration of training hyperparameters given the memory constraints and how specific memory-related hyperparameters affect performance. We investigate how batch size (bs), the number of negatives (ns), and sequence length (sl) affect the NDCG@10 metric of the SASRec model after training. We used a fixed grid of configurations, evaluating approximately $170$ points per dataset, resulting in a total of $997$ runs across all datasets. While the grid does not provide a fully uniform coverage of the hyperparameter space, the analysis can still be interpreted under the assumption of reasonably even exploration (Scarlett et al., 2017; Srinivas et al., 2010).
@@ 5.4.1. Comparison Sampling Strategies
Tables 6, 7 present data to compare metrics and computational resource usage for training with $\mathrm{CCE}^{-}$ using global uniform and popularity-based sampling strategies. In terms of $\mathrm{NDSG@10}$ metric, global uniform sampling consistently outperforms the popularity-based approach across all considered datasets. However, popularity-based sampling shows significant gains in diversity-related metrics for the Gowalla and Beauty datasets. Compared to popularity-based sampling, using global uniform sampling leads to a reduction in $\mathrm{Coverage@10}$ by 185.8% and 63.9 %, and a reduction in $\mathrm{Surprisal@10}$ by 40% and 19.8%. These results align with observations that popularity-weighted sampling enhances exposure to long-tail items, improving recommendation diversity and novelty (Wang et al., 2024).
@@
Another approach to negative sampling is the Scalable Cross-Entropy (SCE) method  (Mezentsev et al., 2024). This method approximates the CE loss by employing a sampling procedure that computes logits only for elements with the largest absolute values, rather than for all items in the catalog. As a result, memory usage for the logit tensor is significantly reduced, and the loss computation is accelerated.
@@ 5.5. Comparison of CCE and $\mathrm{CCE}^{-}$ with BCE
As outlined in the original paper (Kang and McAuley, 2018), the SASRec model can be trained using the Binary Cross-Entropy (BCE) loss. To evaluate the effectiveness of our proposed methods, we conducted training SASRec with the BCE loss.
@@ 5.6. Analysis of efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec
In addition to the SASRec architecture, we study the efficiency of CCE and $\mathrm{CCE}^{-}$ for Bert4Rec  (Sun et al., 2019).